# RF-3DGS: Wireless Channel Modeling with Radio Radiance Field and 3D Gaussian Splatting

### v1

@@ II-D Spatial Spectrum
Space Alternating Generalized Expectation maximization (SAGE) algorithm [47] is an iterative technique for parameter estimation, particularly effective in array signal processing [48, 49, 50, 51]. However, its computational complexity escalates with the number of MPCs, making it less suitable for generating a complete spatial spectrum. Consequently, in our study, SAGE was only employed for extracting MPCs from field measurements.
II-E Conclusion
Based on the above discussion, the following requirements for radio spectra are essential for the success of radiance field reconstruction methods:
  • The spectra should convey maximal geometric information.
  • The objects within the spectra should remain consistent across different views.
  • The spectral intensity should follow transformations that minimize the impact of distance-dependent decay.
Figure 5: RF-3DGS pipeline. (a) The extracted geometry representation from the visual data, consisting of millions of Gaussians, where the Gaussians are visualized by their contour surfaces. (b) The reconstructed radio radiance field, where the objects are illuminated by RF waves. (c) An example of the CSI-encoded spectrum and the corresponding decoded CSI spectra from the respective channels.
III Related Works
Ray tracing: Originally developed for computer graphics, ray tracing has long been adapted for radio propagation modeling [52, 53]. It models radio waves as rays to simulate complex RF interactions such as reflection, diffraction, and scattering [10, 11, 12, 13, 14]. Although ray tracing is suitable for a wide range of environments, it is limited by high computational complexity, stringent requirements for environmental information, and an inflexible pipeline that is difficult to calibrate.
NeRF2 reconstructs a “squared” radio radiance field that considers both “camera” and “light source” locations as inputs. However, its spectra, generated using CBF at 915 MHz with linear scale transformation and per-view normalization, suffer from strong interference, inconsistencies across views, lack of geometric information, and distance-dependent decay, all of which conflict with the requirements discussed earlier. To maintain effectiveness, NeRF2 requires a sample density 2000$\times$ higher than in this work, negating the advantage of sparse sampling that NeRF offers. RayProNet [58] integrates environmental information as input and uses an explicit ”light probe” as the radio radiance field voxel. However, RayProNet is limited to producing only 2D path loss maps, without offering more detailed information.
IV Proposed RF-3DGS Framework
As previously mentioned, although 3DGS has set a new benchmark in optical radiance field reconstruction, two critical challenges must be addressed to adapt it for the RF domain. The first challenge arises from the differences between typical novel view synthesis tasks and wireless communication. In most optical radiance field reconstruction scenarios, scenes are assumed to have static lighting with arbitrary camera poses, and the only point of interest is the photograph. Consequently, the last-bounce-only volume rendering method is well-suited to this context, as it records how the static radiance arrives at the camera from the last interacted object, disregarding prior propagation.
In wireless communication, the situation is more complex. Typically, we also assume a fixed Tx and a moving Rx, with the “picture” being the AoA-gain spatial spectrum. Although this spectrum is valuable, massive MIMO systems often require more detailed CSI on the Tx side, where the larger size array is typically used to achieve more precise beamforming and serve multiple Rxs. Additionally, other CSI features such as delay, phase, and polarization are also vital for wireless communication systems. In RF-3DGS, we address this challenge by developing CSI-encoded SH functions to represent these essential path characteristics.
The second challenge is the difficulty of obtaining sufficient geometric information from the radio spatial spectrum, as discussed in Sec. II. Although massive MIMO systems operating in mmWave bands mitigate this issue to some extent, finding an appropriate method to accurately capture scene geometry remains critical. To address this, RF-3DGS proposes a two-stage fusion training strategy that integrates visual data and RF data for enhanced training, along with multiple array processing techniques, thereby ensuring the effectiveness of radiance field reconstruction.
IV-A RF-3DGS Overview
Building upon the foundations of 3DGS, RF-3DGS adopts a similar representation structure and rasterization rendering pipeline. As illustrated in Fig. 5, RF-3DGS uses visual images and radio spatial spectra as training targets, with the corresponding camera/array poses as input to render the predicted spectra, applying gradient descent to refine the representations.
For geometry representation, RF-3DGS utilizes 3D Gaussian distributions, each with learnable parameters, including density values $\alpha$, positions (mean $(x,y,z)$), and shapes. The shape can be loosely visualized as an “ellipsoid” controlled by a quaternion $\mathbf{q}$ for rotation and a 3D scaling vector $\mathbf{s}\in\mathbb{R}^{3}$, equivalent to the $3\times 3$ covariance matrix of the 3D Gaussian. During the adaptive training process, these quantities—density, positions, and shapes—are optimized to accurately represent the ground truth scene geometry. For view-dependent radiance representation, RF-3DGS assigns CSI-encoded SH functions to each Gaussian, encapsulating the CSI of paths departing from the Gaussian in all directions.
These representations define a field equation that can be directly integrated without relying on complex and lossy sampling strategies. Instead of integrating the rays one by one, RF-3DGS, inheriting from 3DGS, utilizes an efficient rasterization rendering pipeline, which is well-suited to GPU computational structures. This rasterization process can be interpreted as splatting the “ellipsoids” onto the image plane, forming overlapping 2D ellipses that constitute the final image. Consequently, the loss function is defined as the L1 loss plus a differential Structural Similarity Index Measure (SSIM) loss term between the target image and the rendered image. The gradient descent then backpropagates from this loss through the differentiable rasterization pipeline to the learnable parameters of each Gaussian that is visible from the current camera perspective. Further details on how this pipeline handles view splitting, GPU memory management, and rasterization acceleration can be found in [22].
Despite the specifics of the rasterization process, the rendering process for each ray remains equivalent to the continuous volume rendering formula, as shown in Eqn. (1). This implies that the requirements for the spectra remain consistent with those proposed in Sec. II-E.
IV-B CSI-encoded SH Function
To begin, it is essential to understand how SH functions work. This approach employs a linear combination of simple orthogonal basis SH functions to represent a complex spherical function. The number of basis functions is determined by the degree; the higher the degree, the more basis functions are available, allowing for the representation of more complex functions. This means SH functions are most effective if the represented spherical function exhibits continuity in the angle domain. When this continuity is absent, a significantly higher degree of SH functions is required, leading to impractical complexity.
Returning to our discussion, while the last-bounce-only volume rendering may initially seem disconnected from the representation of other CSI features, such as delay and AoD, a more detailed analysis reveals a significant correlation between these features and the path geometry when examined across different paths.
Consider, for instance, a cluster of MPCs reflected from a desk into a fixed Rx array. These MPCs exhibit not only continuous channel gains, as shown in the AoA-gain spectrum (Fig. 1), but also continuous delay and AoD values due to their closely located paths. This correlation extends to MPCs originating from the same Gaussian and arriving at nearby camera positions. The generated CSI spectra (Fig. 5) validate this assumption. These two types of continuity—continuity from one Gaussian to adjacent camera positions and from different AoAs to one camera position—suggest that they can be effectively represented using the SH functions assigned to each distinct Gaussian.
Thus, along with the AoA-gain spectrum, the other CSI features can be encoded using parallel channels, much like how visual images encode RGB values. In this study, we focus on encoding AoD and delay, both of which are critical in wireless communication and sensing. Specifically, for AoD, we decompose it into two channels: one encoding the azimuth angle (0-360) and the other encoding the zenith angle (0-180). Given that the significance of each path is primarily determined by its channel gain, the AoD values are scaled by a gain coefficient, which equals the normalized channel gain within the range of (0,1). The subsequent experimental results demonstrate that RF-3DGS can effectively capture CSI spectra and accurately predict such spectra at test locations.
IV-C Two-stage Fusion Training
From Fig. 4, even the ideal spectrum can only be considered as low-quality “photos” with other spectra offering even less geometric information. Thus, in RF-3DGS, we assume access to both the radio spatial spectra and visual data, which serves as a low-cost supplement to geometric information. The proposed two-stage fusion training process is designed to extract geometric information from the visual data to enhance the radio radiance field representation.
Figure 6: Extrapolation ability of the proposed two-stage fusion training. Both spectra are rendered using the same 300x200 pinhole camera model. The ideal spectrum, serving as the training target, is limited by the maximum MPCs our RTX A6000 48GB GPU can handle and the minimum array pattern lobe width needed for spectrum continuity.
Before beginning the training, if the visual data consists of photos or video without camera intrinsic and extrinsic parameters, an initialization process is required to estimate these parameters and generate initial point clouds using SfM. Otherwise, this process can be skipped, allowing training to start from random point clouds.
Once initialization is complete, the first stage of training begins with a warm-up phase. During this phase, the Gaussians are trained using resolution-reduced images, enabling them to quickly capture global geometry rather than getting stuck on optimizing fine details. The image resolution is doubled twice, after 250 and 500 iterations, until the original resolution is reached.
Following the warm-up, a periodic densification process refines the Gaussians every 100 iterations. In regions where the number of Gaussians is insufficient to represent intricate 3D geometry, the densification process splits large Gaussians and clones small Gaussians to create more learnable “ellipsoids” in those regions. Gaussians with very small density values or those with excessively large footprints in the views are removed to prevent the ”floater” problem, a common issue in radiance field reconstruction methods.
To manage the quantity of Gaussians, their $\alpha$ values are reset towards zero every 3,000 iterations. This reset allows key Gaussians to quickly regain their original density values during optimization, while abnormal or unnecessary Gaussians, whose density values increase slowly, are removed in subsequent densification steps.
After establishing a well-trained geometry representation, the second stage begins. In this stage, RF-3DGS uses the collected radio spectra to train the CSI-encoded SH functions and the density of each Gaussian, while freezing the positions $(x,y,z)$, rotation $\mathbf{q}$, and scaling $\mathbf{s}$. After optimization, the new CSI-encoded SH functions can be interpreted as the objects being re-illuminated by a fixed Tx radio, with the RGB color channels corresponding to different CSI features.
The advantage of the proposed two-stage fusion training is not only that it requires low-cost visual data as a supplement, but also exhibits extrapolation capabilities, as shown in Fig. 6. This capability arises from the accurate visual geometric information, which provides the actual radiance source location, allowing vague or interfered radio spectra to be mapped to the correct locations rather than being assigned to vague floaters, as seen in NeRF2.
V Dataset, Result, and Performance Evaluation
Dataset: In this paper, we constructed a digital twin framework to evaluate the effectiveness of RF-3DGS, comprising comprehensive field measurements in a $14\times 15$ m2 lobby and a digital replica of the same space with a full radio simulation pipeline. The 60 GHz measurement data [59] was provided by the National Institute of Standards and Technology (NIST), USA. For the performance experiments presented in this section, we mainly utilize the simulation-based dataset. The simulator used is Sionna [60], an open-source communication simulator developed by NVIDIA, which is equipped with the Sionna RT [61] ray tracer and supports GPU acceleration.
As shown in Fig. 7, our simulation playground is an accurate 3D replicated model of the NIST lobby with a fixed Tx, and the radio spectra are captured from a set of sparse random Rx positions distributed throughout the lobby. Since the setup is designed to simulate a fully diffuse environment at 60 GHz, the number of paths between each Tx-Rx pair often exceeds 300,000, requiring the array size to be limited to around $10\times 10$. More details on the simulation configuration and settings can be found in our open-source code.
@@ V Dataset, Result, and Performance Evaluation
Performance of Radiance Field Reconstruction: In this part, we focus on evaluating the radiance field reconstruction performance. The dataset, consisting of 800 samples, splits into fixed 160 test samples and 640 training samples. Furthermore, the training samples were randomly selected in proportions of 20%, 40%, 60%, 80%, and 100%.
@@
We compared our method, RF-3DGS, with the Conditional GAN (CGAN) [62], representing black-box neural network approaches, and the SOTA NeRF2. The results are presented in Fig. 8.
@@
In general, RF-3DGS demonstrates significantly better LPIPS scores, requires significantly shorter training times, and offers rendering speeds faster than LTE frame rates. Additionally, RF-3DGS achieves high reconstruction quality with very sparse training samples, with rendering quality substantially degrading only when the sample count drops below approximately 20, depending on the uniformity of the sample distribution. For evaluating other parameters, we utilized the full training set to analyze their impact.
@@
Insufficiency of Statistical Metrics: Another crucial point is that only high-level metrics, such as LPIPS (lower is better), can effectively evaluate our spectra. Traditional statistical metrics, such as the Peak Signal-to-Noise Ratio (PSNR) and SSIM, fall short. This issue has been widely discussed in the CV domain [33, 63]. As illustrated in Fig. 9, even though the spectrum predicted by our method is visibly of higher quality, the PSNR and SSIM yield similar values for both spectra, while only LPIPS reflects the difference accurately.
The primary reason for this metric failure is that our scenario differs significantly from other visual task comparisons. In typical visual tasks, like evaluating compression or super-resolution, outputs are usually close to the perfect ground truth images, resulting in high metric scores (PSNR $>$ 35 dB, SSIM $>$ 0.9, LPIPS $<$ 0.05), where errors are mostly pixel-wise.
Figure 9: Limitation of Statistical Metrics. Statistical metrics are significantly affected by noisy errors, making them insufficient for capturing and emphasizing critical errors such as the loss of geometric structure in the NeRF 2 spectra. High-level metrics, such as LPIPS, are more appropriate for evaluating similarity in our task.
@@
From the comparison, it is clear that the CBF and TCBF spectra impair both reconstruction ability and quality, evident from their poor performance across all metrics. The degradation in reconstruction quality is even more pronounced, with PSNR dropping to 5 dB. This suggests that for CBF and TCBF spectra, strong interference, lack of geometric information, and distance-dependent decay severely limit the effectiveness of radiance field reconstruction in large environments.
In contrast, the MVDR spectrum group exhibits much higher performance in both testing scenarios. More interestingly, in homogeneous spectrum testing, the MVDR spectrum group even results in higher PSNR and SSIM compared to the ideal spectrum group, which may appear counter-intuitive given the ideal spectrum’s higher quality and lower interference. However, as shown in Fig. 6, our method has an extrapolation ability that provides more details than the ideal spectrum. This extrapolation results in lower performance in statistical metrics like PSNR and SSIM but yields similar evaluations from LPIPS, further validating the effectiveness of the LPIPS metric in our tasks.
In summary, these comparisons demonstrate that RF-3DGS achieves significantly higher radio radiance field reconstruction quality, faster training and rendering speed compared to other methods. Additionally, RF-3DGS requires only tens of samples to reconstruct the radiance field across a large lobby while providing an explicit geometric representation. These features and high performance underscore the potential of RF-3DGS in upcoming 6G network applications.
Application in Wireless Communication: To evaluate the effectiveness of RF-3DGS in wireless communication, we consider an application scenario where the Rx-side AoA-gain spectra are used to guide the Rx array in performing angle-domain CBF for signal reception. In this context, the AoA of the maximum-gain path is primary concern. The following results illustrate the angle deviation of RF-3DGS with different types of spectra. As shown, although the CBF and TCBF spectra fail to accurately represent the scene geometry, they can still support simple beamforming tasks, though with some performance loss.
@@
CSI-spectrum Prediction: Another key feature of RF-3DGS is its capability to encode comprehensive CSI. In this paper, we tested its capability to represent the AoD and delay, which are critical in wireless communication.
As shown in Fig. 12, the results indicate the ability of RF-3DGS to reconstruct Rx-side AoA-AoD spectra and AoA-delay spectra. We also present examples of the decoded CSI spectra in Fig. 5. For the delay channel, we use normalized delay values, measured from the first arrival of MPCs, which is more practical and helps mitigate distance-dependent decay. The example demonstrates that the delay increases from the near reflection point to the far end, similar to a depth image. In the AoD-azimuth channel, although the angles vary according to physical laws, a sharp edge corresponding to the transition between 0 and 360 is visible. A more appropriate approach would involve using a 3D unit vector to represent such AoD information; however, in this paper, we use these two angles for a more intuitive demonstration.
@@ VI Field Study: A Digital Radio Twin
In previous experiments, our testing relied on simulation datasets, raising concerns about potential discrepancies between our digital twin and its physical counterpart. In this section, we demonstrate the fidelity of our digital radio twin framework, which includes both environment geometry and a comprehensive radio radiance field, and highlight its support for advanced applications like ISAC.
@@
We also conducted field measurements at the same Rx locations using a synchronized Tx-Rx system, as depicted in Fig. 7. This system generates high-definition, precise Channel Impulse Response (CIR) between each Tx and Rx array element pair, which are processed by the SAGE algorithm to estimate key MPCs, including path loss, delay, AoD, and AoA. More details about this system are available in [51, 59]. Using these MPCs, we generated an MPC spectrum similar to our ideal spectra, which was then used to validate and calibrate our digital twin. The lower image of Fig. 13 shows the comparison results of direct validation and after calibration.
@@
This scenario is common when real systems have imperfections. In these cases, the advantages of RF-3DGS, such as rapid training and extrapolation ability, become particularly valuable. With just 3 minutes of re-training, we can produce a validated radio radiance field representation with high testing metrics, as shown in Fig. 14. The extrapolation capability also allows us to map imperfect measurements onto the correct positions.
Figure 13: Comparison of (a) digital twin and (b) physical twin measurements [59].
@@
An ISAC Scenario Demo: Considering RF-3DGS’s remarkable ability to reconstruct the radio radiance field and model the environment with real-time wireless communication feedback, its potential in ISAC applications [29] becomes evident. Fig. 15 illustrates an ISAC scenario. On the left, the spectra rendered using the well-trained RF-3DGS representation serve as prior knowledge. Then, a person enters the scene. Traditional methods might struggle to capture and render such dynamic changes in real-time [64, 65]. However, with RF-3DGS, the system not only detects the individual’s presence but also quickly generates a highly detailed radio spatial spectrum from the communication feedback, accurately reflecting the person’s shape and location within the scene by comparing the captured spectra with those rendered from the prior representation
This capability is particularly beneficial in environments where timely and accurate sensing is crucial, such as in safety monitoring or automated navigation systems. RF-3DGS enables the simultaneous acquisition of communication signals and precise reconstruction of the surrounding environment’s geometry in an explicit manner.
Figure 15: ISAC scenario demonstration. The left figures depict the well-trained radio radiance field representation, which serves as prior knowledge. The right figures illustrate the real-time scenario, where the wireless channel feedback is compared to the prior to locate the sensing target.

### v2

@@ IV-E Generating Radio Spatial Spectrum
Space Alternating Generalized Expectation maximization (SAGE) algorithm [41] is an iterative technique for parameter estimation, particularly effective in array signal processing [42, 43]. Therefore, in our study, SAGE was used exclusively for extracting MPCs from field measurements, and the extracted MPCs were subsequently used to generate the spectrum. As shown in Fig. 14(b), the SAGE spectrum is generated by an equirectangular camera model and overlaid with a picture from a similar optical camera at same pose.
V Experimental Result and Evaluation
In this section, to evaluate the performance of RF-3DGS, we designed several experiments. First, we directly compared the performance of RRF reconstruction with other methods. Next, to assess its generalizability, we tested how different types of spectra affect the performance of RF-3DGS. To evaluate RF-3DGS’s effectiveness in massive MIMO spatial management, we conducted experiments to test its ability to predict the CBF steering vector. Finally, as mentioned, RF-3DGS reconstructs multi-dimensional CSI, and we tested the accuracy of the reconstructed AoD and delay.
@@ V Experimental Result and Evaluation
Performance of Radio Radiance Field Reconstruction: In this part, we focus on evaluating the radiance field reconstruction performance. The dataset, consisting of 800 samples, splits into fixed 160 test samples and 640 training samples. Furthermore, the training samples were randomly selected in proportions of 20%, 40%, 60%, 80%, and 100%.
A important point is that only high-level metrics, such as LPIPS (lower is better), can effectively evaluate our spectra quality when compared to the ground truth. Traditional statistical metrics, such as the Peak Signal-to-Noise Ratio (PSNR) and SSIM, fall short. This issue has been widely discussed in the CV domain [44, 45]. As illustrated in Fig. 9, even though the spectrum predicted by our method is visibly of higher quality, the PSNR and SSIM yield similar values for both spectra, while only LPIPS reflects the difference accurately.
The main reason for this metric failure is that our scenario differs significantly from other visual task comparisons. In typical visual tasks, like evaluating compression or super-resolution, outputs are usually close to the perfect ground truth images, resulting in high metric scores (PSNR $>$ 35 dB, SSIM $>$ 0.9, LPIPS $<$ 0.05), where errors are mostly pixel-wise.
Figure 9: Limitation of Statistical Metrics.
@@
With LPIPS as the primary metric, we compared our method, RF-3DGS, with the Conditional GAN (CGAN) [46], representing black-box neural network approaches, and the SOTA NeRF2. The results are presented in Fig. 10.
@@
RF-3DGS demonstrates significantly better LPIPS scores, requires orders of magnitude shorter training times, and offers rendering speeds faster than Long Term Evolution (LTE) frame rates. Additionally, RF-3DGS achieves high reconstruction quality with very sparse training samples, with rendering quality substantially degrading only when the sample count drops below approximately 20, depending on the uniformity of the sample distribution. For evaluating other parameters, we utilized the full training set to analyze their impact.
@@
From the comparison, it is clear that the CBF and TCBF spectra impair both reconstruction ability and quality, evident from their poor performance across all metrics. The degradation in reconstruction quality is even more pronounced, with PSNR dropping to 5 dB. This suggests that for CBF and TCBF spectra, strong interference, and lack of geometric information severely limit the effectiveness of radiance field reconstruction in large environments.
In contrast, the MVDR spectrum group exhibits much higher performance than CBF and TCBF in both testing scenarios. More interestingly, in homogeneous spectrum testing, the MVDR spectrum group even results in higher PSNR and SSIM compared to the ideal spectrum group, which may appear counter-intuitive given the ideal spectrum’s higher quality and lower interference. However, as shown in Fig. 6, our method has an extrapolation ability that provides more details than the ideal spectrum. This extrapolation results in lower performance in statistical metrics like PSNR and SSIM but yields similar evaluations from LPIPS, further validating the effectiveness of the LPIPS metric in our tasks.
In summary, these comparisons demonstrate that RF-3DGS achieves significantly higher radio radiance field reconstruction quality, faster training speed, and faster rendering speed compared to other methods. Additionally, RF-3DGS requires only tens of samples to reconstruct the radiance field across a large lobby while providing an explicit geometric representation. These features and high performance underscore the potential of RF-3DGS in future 6G network applications.
Application in MIMO Spatial Management: To evaluate the effectiveness of RF-3DGS in wireless communication, we consider an application scenario where the Rx-side channel gain spectra are used to guide the Rx array in performing angle-domain CBF for signal reception. In this context, the AoA of the maximum-gain path is primary concern. The following results illustrate the angle deviation of RF-3DGS with different types of spectra. As shown in Fig. 12, although the CBF and TCBF spectra fail to accurately represent the scene geometry, they can still support simple beamforming tasks, though with some performance loss.
@@
Spatial-CSI Prediction: Another key feature of RF-3DGS is its capability to encode Spatial-CSI. In this paper, we tested its capability to represent the AoD and delay, which are critical in wireless communication.
As shown in Fig. 13, the results indicate the ability of RF-3DGS to reconstruct Rx-side AoD spectra and delay spectra. We also present examples of the decoded CSI spectra in Fig. 5. For the delay channel, the example demonstrates that the delay increases from the near reflection point to the far end, similar to a depth image. In the AoD-azimuth channel, although the angles vary according to physical laws, a sharp edge corresponding to the transition between 0 and 360 is visible. A more appropriate approach would involve using a 3D unit vector to represent such AoD information; however, in this paper, we use these two angles for a more intuitive demonstration.
@@ VI Field Study: A Digital Radio Twin
In previous experiments, our testing relied on simulation datasets, which raised concerns about potential discrepancies between our digital twin and its physical counterpart. In this section, we demonstrate the fidelity of our digital radio twin framework, which includes both environment geometry and a comprehensive radio radiance field, and highlight its support for advanced applications like ISAC.
Figure 14: (a) digital twin and (b) physical twin measurements comparisons [34].
@@
We also conducted field measurements at the same Rx locations using a synchronized Tx-Rx system, as depicted in Fig. 7. This system generates high-definition, precise Channel Impulse Response (CIR) between each Tx and Rx array element pair, which are processed by the SAGE algorithm to estimate key MPCs, including path loss, delay, AoD, and AoA. More details about this system are available in [43, 34]. Using these MPCs, we generated an MPC spectrum similar to our ideal spectra, as shown in the lower image of Fig. 14, which was then used to validate and calibrate our digital twin.
@@
This scenario is common when real systems have imperfections. In these cases, the advantages of RF-3DGS, such as rapid training and extrapolation ability, become particularly valuable. With just 3 minutes of re-training, we can produce a validated radio radiance field representation with high testing metrics, as shown in Fig. 15.
@@
An ISAC Scenario Demo with Digital Radio Twin: Considering RF-3DGS’s exceptional ability to reconstruct the radio radiance field and model the environment with real-time wireless communication feedback, its potential for ISAC applications [21] is evident. Fig. 16 illustrates an ISAC scenario within digital radio twin, which consists of a normal digital twin framework and a well-trained RF-3DGS representation. On the left of Fig. 16, the ability of RF-3DGS to provide real-time radio spatial spectra at arbitrary positions serves as prior knowledge of the radio scattering background. When a sensing target enters the scene as shown on the right of Fig. 16, optical monitors detect the target’s entry. However, visual data alone can only provide vague information about the target, limiting its utility for precise sensing. Such vague information can be used by the digital twin to notice the connected Rxs the relative location of the sensing target within their FoV, as highlighted by the red boxes. With both the real-time radio scattering background and the target’s potential direction, wireless communication feedback can be processed more effectively to extract detailed information about the sensing target. In comparison, traditional methods often struggle to distinguish and render such dynamic changes from wireless communication feedback in real time[47, 48].
Figure 16: ISAC scenario demonstration.
This capability is particularly advantageous in scenarios where timely and accurate sensing is critical, such as safety monitoring or automated navigation systems. RF-3DGS facilitates the simultaneous acquisition of communication signals and the precise reconstruction of the surrounding environment’s geometry in an explicit and efficient manner.